### v1

# How Well do LLMs Compress Their Own Chain-of-Thought? A Token Complexity Approach

Abstract
Chain-of-thought prompting has emerged as a powerful technique for enabling large language models (LLMs) to solve complex reasoning tasks. However, these reasoning chains can be verbose, raising concerns about efficiency. In response, recent works have sought to decrease response lengths through simple prompting strategies (e.g. ‘be concise’). In this work, we conduct the first systematic study of the relationship between reasoning length and model performance across a diverse range of compression instructions (e.g. ‘use 10 words or less’ or ’remove all punctuation’). In doing so, we discover a universal tradeoff between reasoning length and accuracy that persists across even very distinct reasoning chains. We demonstrate that this tradeoff emerges from a sharp threshold behavior at the question level: each task has an intrinsic ‘token complexity’ – a minimal number of tokens required for successful problem-solving. We show how token complexity enables us to compute information-theoretic limits on the accuracy-compression tradeoff, and find that prompt-based compression strategies operate far from these theoretical limits. This suggests there may be significant room for improvement and our framework provides a benchmark to help researchers evaluate progress in reasoning efficiency. Our work also highlights the importance of adaptive compression – giving shorter responses for easier questions – and we show that token complexity is a useful tool for measuring this capability.111 Our code is available at https://github.com/Compressed-CoT/compressed-cot. Our data is available at https://huggingface.co/datasets/cot-compression/TokenComplexity
How Well do LLMs Compress Their Own Chain-of-Thought?
A Token Complexity Approach
Ayeong Lee Columbia Business School al3393@columbia.edu                        Ethan Che Columbia Business School ewc2119@columbia.edu                        Tianyi Peng Columbia Business School tp2845@columbia.edu
1 Introduction
Recent advancements in large language models (LLMs)—including GPT-o1 and DeepSeek R1—alongside broader AI agent development, have showcased impressive reasoning capabilities, hinting at a future where complex problem-solving and decision-making can be automated. However, this rapid progress also introduces a significant challenge: the computational cost of reasoning is projected to increase substantially as these models are deployed in real-world applications. This growing inference cost highlights the necessity for efficient reasoning strategies, motivating our research into reducing the computational burden of LLM inference while maintaining high performance.
A pivotal technique for enhancing LLM reasoning has been chain-of-thought (CoT) prompting, which encourages models to generate intermediate reasoning steps before arriving at a final answer Wei et al. (2022); Kojima et al. (2022). While effective, these reasoning chains often involve lengthy intermediate computations, significantly increasing inference costs when deployed at scale Yu et al. (2024).
Although CoT-generated outputs are typically verbose, containing redundant Chiang and Lee (2024) or sometimes irrelevant information Wang et al. (2023), it remains unclear how best to compress the chain-of-thought content for effective problem-solving. Prior work has observed that asking the LLM to ‘be concise’ or ‘use at most 100 words’ can reduce response length while incurring a range of degradation to accuracy Jin et al. (2024); Renze and Guven (2024); Han et al. (2024). Yet so far, there has not been a comprehensive exploration of the effectiveness of different compression strategies. For example, should language models use fewer reasoning steps instead of using fewer words or characters? Should they be removing unnecessary punctuation or filler words (e.g. ‘therefore’, ‘so’). Should they reason in alternative languages that may be more token-efficient?
Prompt Id Prompt Sample Output under Claude 3.5 Sonnet
NoCoT Only give the final answer. Answer: (G)
DefaultCoT Think step-by-step. Let me solve this step by step: 1) Z18 (…)
BeConcise (…), but be concise. Let’s solve this step by step: 1) Z18 (…)
BulletPoints (…) only use bullet points. - Z18 is the group of integers modulo 18 (…)
OnlyNumbers (…) only use numbers or equations. |Z18| = 18. Answer: (G)
NoSpaces (…) do not use any spaces or line breaks. TheorderofagroupZnisthesamenumbern,sothe (…)
NoProperGrammar (…) do not use proper grammar. lemme help u with this z18 is group
AbbreviateWords (…) abbreviate words as much as possible. Slvng fr rdr f Z18: Z18 = grp f ntgrs mdle (…)
WordLimit(k) (…) use at most $k$ words. $k\in\{1,...,100\}$ Order of Z18 is eighteen. Answer: (G)
CharLimit(k) (…) use at most $k$ letters. $k\in\{1,...,500\}$ The order of a group is the number of elements in (…)
TokenLimit(k) (…) use at most $k$ tokens. $k\in\{1,...,500\}$ The order of Z18 is 18, as it contains elements (…)
StepLimit(k) (…) use at most $k$ steps. $k\in\{1,...,5\}$ Step 1: The order of a group Zn is simply n, so (…)
ChineseCoT (…) Respond in Chinese {CJK} UTF8gbsn让我帮你解答这个问题. Z18 是-个循环群 (…)
ChineseCoT(k) (…) Use at most $k$ Chinese characters. $k\in\{1,...,500\}$CJK} UTF8gbsn答案是18. Answer: (G).
Table 1: The Chain-of-Thought prompts we consider. The right column gives an example of the chain-of-thought of Claude 3.5 Sonnet on a sample problem in MMLU-Pro Math.
The main contribution of our work is to provide the first systematic study of the trade-off between reasoning length and performance across different prompt-based compression strategies, including prior strategies such as ‘be concise’ as well as alternative approaches such as ‘only use bullet points’ or ‘use at most 50 Chinese characters’. In total, we evaluate 31 prompts for six LLMs on three standard reasoning datasets. Remarkably, although these prompting strategies produce different chains of thought, their trade-offs between response length and accuracy lie on a universal trade-off curve. In other words, all prompts are equally "good" as extremes on this curve. What primarily affects accuracy is the length of the chain of thought, far more than changes in its composition.
Our second contribution is a novel empirical observation: the performance of reasoning tasks exhibits a sharp threshold dependence on reasoning length at the question level. By evaluating multiple prompts for each question, we demonstrate that most questions have a well-defined ‘token complexity’—a minimum number of tokens required to successfully solve the question. This token complexity is universal across prompting strategies. We estimate token complexities across various benchmarks and find that: (i) Token complexity alone can predict the performance of CoT prompting strategies with 95% accuracy. (ii) It serves as a robust measure of reasoning task difficulty, enabling us to investigate whether LLMs reason adaptively—using shorter chains-of-thought for easier questions.
These results raise the question of whether the trade-off curve between response-length and accuracy induced by these prompting strategies are close or far from optimal. Viewing these strategies as a form of ‘lossy compression’, we borrow tools from rate-distortion theory to characterize the optimal accuracy-compression trade-off. In doing so, we find that prompt-based strategies are far from the theoretical limits, especially on harder datasets. In particular, our analysis highlights the importance of adaptive compression and shows that while LLMs tend to use shorter reasoning chains for easier problems, there is significant room for improvement. In total, this work provides a set of empirical tools that can help researchers benchmark new methodologies for improving the efficiency of LLM reasoning.
1.1 Related Work
Recent research has begun to gain traction in exploring the trade-off between response length and accuracy in LLMs. Studies such as Renze and Guven (2024); Jin et al. (2024); Nayab et al. (2024); Han et al. (2024) have employed specific prompting strategies to constrain response length and assess the associated impact on accuracy and performance. Additionally, several works have highlighted the redundancy inherent in CoT prompting Chiang and Lee (2024); Wu et al. (2025) and emphasized the benefits of concise reasoning. Other approaches have focused on fine-tuning strategies to adapt LLMs for generating more succinct reasoning Kang et al. (2024); Yu et al. (2024).
Our work advances this growing body of literature by making three key contributions: 1) We conducted a systematic evaluation of a rich set of prompts designed to reduce the length of CoT reasoning while maintaining accuracy; 2) We find that accurate answers are only achieved when the output length exceeds a certain threshold, which is intrinsic to the problem and independent of the CoT format. We formalize this concept as the token complexity of a problem; 3) We derive theoretical limits on the length-accuracy tradeoff, providing a framework for researchers to benchmark new methodologies aimed at compressing chain-of-thought reasoning effectively.
2 Experiments
Our evaluation encompasses the following LLMs: GPT-4o Hurst et al. (2024), GPT-4o-mini Hurst et al. (2024), Claude 3.5 Sonnet Anthropic (2024), Claude 3.5 Haiku Anthropic (2024), and Llama 3.3 70B Instruct Dubey et al. (2024). We evaluate these models on three standard math reasoning datasets: MATH-500 Lightman et al. (2023), a random 500 problem subset of GSM8K (GSM8K, Cobbe et al. (2021)), and a random 500 problem subset of MMLU-Pro Math problems (MMLU-Pro Math, Wang et al. (2024)).
For each LLM and dataset, we test 31 prompts designed to induce shorter response lengths, detailed in Table 1. These prompts include ones considered in prior literature: ‘be concise’ Renze and Guven (2024), ‘use $k$ words or less’ Jin et al. (2024); Nayab et al. (2024), ‘use $k$ tokens or less’ Han et al. (2024), but include additional curated ones to assess the impact of alternative compression strategies. For each prompt, we assess performance with two metrics: (1) accuracy, the fraction of questions solved correctly, and (2) average token length, the average number of output tokens produced by the LLM in their response across questions in the dataset.
Model Prompt Accuracy Token Count GPT-4o NoCoT 42% 10 DefaultCoT 81% 586 BeConcise 80% 415 BulletPoints 75% 185 OnlyNumbers 79% 222 NoSpaces 79% 249 NoProperGrammar 79% 189 AbbreviateWords 74% 276 ChineseCoT 75% 326 Claude 3.5 Sonnet NoCoT 48% 8 DefaultCoT 81% 324 BeConcise 78% 227 BulletPoints 76% 167 OnlyNumbers 77% 130 NoSpaces 77% 188 NoProperGrammar 77% 173 AbbreviateWords 77% 220 ChineseCoT 76% 311 Llama-3.3-70B NoCoT 43% 20 DefaultCoT 74% 552 BeConcise 77% 443 BulletPoints 70% 194 OnlyNumbers 63% 304 NoSpaces 67% 330 NoProperGrammar 68% 217 AbbreviateWords 70% 327 ChineseCoT 74% 363
Table 2: Comparison of Accuracy and Average Token Length of chain-of-thought prompts on MMLU-Pro Math.
Figure 1: For each of the 31 CoT prompts we consider (see legend above), we report average token length vs accuracy for GPT-4o and Claude 3.5 Sonnet on MMLU-Pro Math and GPT-4o-mini on GSM8K tasks. Despite differences in the chain-of-thought, many of them live on a universal tradeoff curve. We report upper bounds for optimal accuracy under a given average token budget (see section 4). See Appendix E for more models and benchmarks.
2.1 Benchmark Results
By considering multiple diverse prompts, we are able to induce chains-of-thought along a spectrum of response lengths and reasoning performance, with NoCoT (no chain-of-thought) using the fewest tokens with the lowest accuracy and DefaultCoT (i.e. standard chain-of-though-prompting ‘think step-by-step’) using the most tokens and generally having the highest benchmark performance. The dataset is available at https://huggingface.co/datasets/cot-compression/TokenComplexity. In Table 2, we focus on a subset of chain-of-thought prompts and report the accuracy and average token count for MMLU-Pro Math across several LLMs. We highlight the following observations, which also hold for other datasets (see Appendix A).
  • There is potential to achieve significant length reduction (e.g., up to 60%) compared to DefaultCoT without sacrificing much accuracy.
  • BeConcise Renze and Guven (2024) consistently reduces token length without significantly hurting performance.
  • Yet, there are other prompts such as OnlyNumbers or NoProperGrammar that preserve a similar accuracy as BeConcise but induce substantially shorter chains-of-thought (a $\approx 50\%$ reduction for GPT-4o). The strong performance of these prompts suggests that LLMs do not necessarily require proper English to conduct chain-of-thought reasoning effectively.
  • There is no universally dominant prompt; while OnlyNumbers and NoProperGrammar generally work well for GPT-4o and Claude 3.5 Sonnet, they are less performant for LLaMA 3.3 70B.
  • While LLMs do not always follow the prompt exactly, the prompt still induces large reductions in the response length (as also observed in  Han et al. (2024)).
As researchers (e.g. Kang et al. (2024)) propose new methodologies for improving the accuracy-length trade-off, our experimental results provide a simple yet effective baseline.
2.2 Universal Trade-off between Reasoning Length and Accuracy
Figure 2: Evidence of the token complexity hypothesis. (Left) Performance of Claude 3.5 Sonnet on a sample question in MATH-500 exhibits a threshold behavior. With the exception of 2 prompts, all the prompts that use more tokens than the threshold get the answer correct, the rest do not. The red dotted line indicates the estimated token complexity $\tau_{i}^{\pi}$ from the results. (Right) Actual accuracy on MATH-500 vs predicted accuracy from the token complexity hypothesis. Token complexity is highly predictive of actual accuracy.
The observation that the token length of DefaultCoT can be substantially improved upon without much degradation to accuracy motivates a natural question: which prompts exhibit the best tradeoff between response length and accuracy?
To study this question, we plot the average token-length and accuracy of all 31 prompts in Figure 1 for the MMLU-Pro Math benchmark (results for other benchmarks are in Appendix E). Remarkably, we see that almost all the prompts we consider lie on a universal trade-off curve between response length and accuracy. This suggests that regardless of whether the chain-of-thought is formatted in bullet points, without spaces, using only numbers, or even in Chinese, ultimately it is the length of the chain-of-thought that matters most. This result also holds for the Wordlimit(k), Charlimit(k), TokenLimit(k), StepLimit(k), and ChineseCoT(k) prompts, which ask the LLM to limit the response to be at most $k$ words, letters, tokens, reasoning steps, or Chinese characters. One would expect that each of these strategies would result in separate tradeoff curves, but surprisingly all of these strategies result in a near-identical trade-off between response length and accuracy. This suggests that there is not much room for improvement simply by changing the composition of the chains-of-thought.
This universal trade-off curve suggests that the length of the chain-of-thought is the predominant factor that influences reasoning performance. We caveat this observation acknowledging that this universal trade-off should only hold for reasonably informative chains-of-thought, i.e. we would expect that pure white-space would perform worse. We also see that adherence to the universal trade-off curve is better for more capable models (i.e. GPT-4o and Claude-3.5-Sonnet) on easier benchmarks (i.e. GSM8K). For less capable models such as LLaMA 3.3 70B on harder datasets (e.g. MATH-500), there are more prompts which are below the trade-off curve.
3 The Token Complexity Hypothesis
The results from the previous section highlight the importance of response length on reasoning performance. To investigate this relationship further, we use our dataset to study reasoning performance at question-level granularity. Our extensive coverage of response lengths allows us to observe that reasoning performance at the question-level exhibits a sharp threshold-like behavior: across all prompts, the LLM correctly solves the question if and only if the response length is above a certain threshold. We refer to this threshold as the token complexity of the problem.
To illustrate, the left panel of Figure 2 displays the performance of all 31 prompts for Claude 3.5 Sonnet on a sample question in the MATH-500 dataset. Despite the diversity of prompting strategies we consider, we see that response length is highly predictive of correctness: with the exception of 2 prompts, all the prompts which use more than $\approx 53$ output tokens correctly solve the problem, while the prompts that use fewer tokens get the question wrong.
To formalize this behavior, we first introduce some notation. Given a dataset of $i=1,...,n$ questions ($n=500$ for the benchmarks we consider), let $P_{k}$ denote a chain-of-thought prompt and let $X_{i,k}$ denote a chain-of-thought produced by an LLM $\pi$ for question $i$ when prompted by $P_{k}$. We let $t(X_{i,k})\in\mathbb{N}$ be the length of $X_{i,k}$ in tokens. We let $a^{\pi}_{i}(X_{i,k})=1$ if $\pi$ gets the answer correct under $X_{i}$ and 00 if not. We now turn to formally outlining the ‘token complexity hypothesis’.
Assumption 1.
(Token complexity hypothesis) For question $i$ and LLM $\pi$, there exists a threshold $\tau^{\pi}_{i}\in\mathbb{N}$, denoted as the token complexity, such that for any prompt $P_{k}$, the LLM gets the answer correct iff the token length $t(X_{i,k})$ is above $\tau^{\pi}_{i}$:
$a^{\pi}_{i}(X_{i,k})=\mathbf{1}\{t(X_{i,k})\geq\tau^{\pi}_{i}\}$
As discussed before, we should think of this hypothesis as holding only for ‘reasonable’ chains-of-thought (e.g. not 100 tokens of pure whitespace). Nonetheless, as we will discuss, we observe this to hold broadly for all the prompts we test. This universality makes the token complexity notion useful because it provides fine-grained notion of problem difficulty that doesn’t depend on a specific prompting strategy: more difficult problems simply require more tokens to solve.
Dataset Model DefaultCoT Token Count BeConcise Token Count $T_{\pi}^{*}(A^{*})$ BeConcise Token Reduction Upper Bound of Token Reduction Math-500 GPT-4o 635 505 172 1.26x 3.69x GPT-4o-mini 611 528 164 1.16x 3.72x Claude-3.5-Sonnet 373 283 105 1.32x 3.56x Claude-3.5-Haiku 373 287 143 1.30x 2.61x Llama-3.3-70B 549 475 93 1.16x 5.88x GSM8K GPT-4o 266 190 24 1.40x 10.90x GPT-4o-mini 292 216 35 1.35x 8.29x Claude-3.5-Sonnet 200 136 18 1.47x 11.16x Claude-3.5-Haiku 212 169 42 1.25x 5.00x Llama-3.3-70B 195 148 27 1.32x 7.24x MMLU-Pro GPT-4o 586 415 121 1.41x 4.83x Math GPT-4o-mini 506 419 132 1.21x 3.85x Claude-3.5-Sonnet 324 227 59 1.43x 5.50x Claude-3.5-Haiku 296 237 93 1.25x 3.20x Llama-3.3-70B 552 443 75 1.25x 7.36x
Table 3: Comparison of token counts on the MMLU-Pro Math, GSM8K, and Math-500 datasets under DefaultCoT and BeConcise along with $T_{\pi}^{*}(A^{*})$, which is a lower bound on the average token length required to achieve max accuracy. While existing prompt strategies do meaningfully reduce token-length, the lower bound $T_{\pi}^{*}(A^{*})$ illustrates that one can achieve drastically more compression while preserving accuracy.
We proceed to test this hypothesis quantitatively across LLMs $\pi$ and benchmarks. First, we measure to what degree success or failure on a task can be classified based purely on the token-length of the chain-of-thought, i.e. whether the behavior in the left panel of Figure 2 holds broadly. To do so, we use our dataset of $k=1,...,K=31$ chain-of-thought prompts for each question. For $t\in\mathbb{N}$, we define $c^{\pi}_{i}(t)$ to be the classification accuracy under a threshold classifier under $t$:
$c^{\pi}_{i}(t)\equiv\frac{1}{K}\sum_{k=1}^{K}\mathbf{1}\{a_{i}^{\pi}(X_{i,k})=\mathbf{1}\{t(X_{i,k})\geq t\}\}$
Our estimator $\hat{\tau}_{i}^{\pi}$ of token complexity is the optimal threshold-based classifier, and we let $c_{i}^{*}$ be the maximum classification accuracy achieved by $\hat{\tau}^{\pi}_{i}$.
$\hat{\tau}^{\pi}_{i}$ $\equiv\arg\max_{k}c_{i}^{\pi}(t(X_{i,k}))$ (1)
$c_{i}^{*}$ $\equiv\max_{k}c_{i}^{\pi}(t(X_{i,k}))$ (2)
$\bar{c}_{\pi}$ $\equiv\frac{1}{n}\sum\nolimits_{i=1}^{n}c_{i}^{*}$ (3)
We let $\hat{\tau}_{i}^{\pi}=\infty$ if setting the threshold to $t=\infty$ results in better classification accuracy, e.g. if none of the chains-of-thoughts correctly solve the problem in which case $a_{i}^{\pi}(X_{i,k})=0$ for all $k$. For each dataset and model, we report the average classification accuracy $\bar{c}_{\pi}$ under the estimated $\tau^{\pi}_{i}$ in Table 4. Overall, the average classification accuracy is very high, above 90% for most models and benchmarks, verifying that (1) token-length is highly predictive for performance at the question-level and (2) question-level performance exhibits a threshold relationship with token-length. And we see that the threshold classifier achieves higher average classification accuracy for (1) more capable models (i.e. GPT-4o and Claude 3.5 Sonnet) and (2) easier benchmarks (i.e. GSM8K), which mirrors the finding in Section 2.2 concerning adherence to the universal trade-off curve.
Our second empirical validation is to compare the accuracy of the LLM across prompts and benchmarks with the performance predicted under the token complexity hypothesis. We let $\text{Acc}_{\pi}(P_{k})$ denote the accuracy of LLM $\pi$ under prompt $P_{k}$ while $\widehat{\text{Acc}}_{\pi}(P_{k})$ denotes the accuracy predicted under our estimated token-complexities:
$\text{Acc}_{\pi}(P_{k})$ $\equiv\frac{1}{n}\sum\nolimits_{i=1}^{n}a^{\pi}_{i}(X_{i,k})$ (4)
$\widehat{\text{Acc}}_{\pi}(P_{k})$ $\equiv\frac{1}{n}\sum\nolimits_{i=1}^{n}\mathbf{1}\{t(X_{i,k})\geq\hat{\tau}^{\pi}_{i}\}$ (5)
$\text{Err}_{\pi}=\frac{1}{K}$ $\sum\nolimits_{k=1}^{K}\frac{|\text{Acc}_{\pi}(P_{k})-\widehat{\text{Acc}}_{\pi}(P_{k})|}{\text{Acc}_{\pi}(P_{k})}$ (6)
In Table 5, for each LLM and dataset we report $\text{Err}_{\pi}$, the average relative discrepancy between actual accuracy and predicted accuracy. We observe that the token threshold classifier is able to predict overall benchmark performance within 6% error, and the error is even smaller for larger models on easier benchmarks. This illustrates that the token-complexity hypothesis provides a model for reasoning performance that is highly predictive of actual performance, which can be seen visually in the right panel of Figure 2 for Claude 3.5 Sonnet on the MATH-500 dataset. Since predicted accuracy $\widehat{\text{Acc}}_{\pi}(P_{k})$ only depends on the token lengths, this also provides an explanation for why accuracy across the chain-of-thought prompts we consider is highly dependent on average token-length in Figure 1. We explore the implications of this hypothesis in the next section.
4 Theoretical Limits of the Length-Performance Tradeoff
Not only does token-complexity provide an interpretable and accurate model of reasoning task performance, it gives insight into how to improve the token-length efficiency. Viewing these response-reducing CoT prompts as a form of lossy compression, this raises a natural research question: what are the theoretical limits on the optimal tradeoff token-length and accuracy and how far away are existing prompting strategies from this limit? Although there are recent works which propose methodologies for improving this tradeoff, so far there has not been a study of the theoretical limits of compression, which makes it difficult to evaluate these proposed methods.
We develop a framework to empirically compute information-theoretic lower bounds on performance, inspired by rate-distortion theory. First, we define $\bar{t}_{\pi}(P)\equiv\frac{1}{n}\sum_{i=1}^{n}t(X_{i})$ to be the average token-length under CoT prompt $P$ and LLM $\pi$. We define $\alpha^{*}(T)$ to be the optimal performance for an average token budget of $T$ tokens, and $T^{*}(\alpha)$ is the minimum average token length to achieve a target accuracy of $\alpha$:
$\alpha^{*}_{\pi}(T)$ $=\max_{P}\{\text{Acc}_{\pi}(P):\bar{t}_{\pi}(P)\leq T\}$ (7)
$T^{*}_{\pi}(\alpha)$ $=\min_{P}\{\bar{t}_{\pi}(P):\text{Acc}_{\pi}(P)\geq\alpha\}$ (8)
These quantities involve intractable optimizations over CoT prompts $P$. Yet, under the token complexity hypothesis, these optimization problems are greatly simplified, involving only optimization over token-lengths instead of prompts.
In fact, the optimization is a special case of a knapsack problem, and thus gives rise to a closed-form solution. First, define the empirical CDF of the true token complexities $F_{n}(t)\equiv\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\{\tau_{i}^{\pi}\leq t\}$, let $E_{n}(t)\equiv\frac{1}{n}\sum_{i=1}^{n}\tau_{i}^{\pi}\mathbf{1}\{\tau_{i}^{\pi}\leq t\}$, and let $\bar{\tau}_{\pi}=\frac{1}{N}\sum_{i=1}^{n}\tau_{i}^{\pi}\mathbf{1}\{\tau_{i}^{\pi}<\infty\}$ denote the average token complexity among questions that have finite token complexity.
Theorem 1.
Suppose Assumption 1 and suppose that for any token counts $\{t_{i}\}_{i=1}^{n}$ there exists a prompt $P$ such that $t(X_{i})=t_{i}$. Then,
$\alpha^{*}_{\pi}(T)$ $=\frac{1}{n}\sum\nolimits_{i=1}^{n}\mathbf{1}\{\tau^{\pi}_{i}\leq t_{T}\}$ (9)
$T^{*}_{\pi}(\alpha)$ $=\frac{1}{n}\sum\nolimits_{i=1}^{n}\tau^{\pi}_{i}\mathbf{1}\{\tau^{\pi}_{i}\leq q_{\alpha}\}$ (10)
where $t_{T}\equiv\inf\{t\in\mathbb{R}:E_{n}(t)=t\}$ and $t_{T}=\infty$ if $t>\bar{\tau}^{\pi}$ and $q_{\alpha}=\sup\{t\in\mathbb{R}:F_{n}(t)\leq\alpha\}$ is the $\alpha$-quantile of the empirical distribution.
Intuitively, under Assumption 1 the optimal strategy is to only use the minimal number of tokens required to solve the problem, $\tau_{i}^{\pi}$. Under a limited token budget, it is optimal to sort questions by token complexity and solve the questions with shortest token complexity until the budget is filled. Note that this structure is due to the fact that each question is weighted identically, if certain questions had a higher weight than others (e.g. harder questions are more valuable) then the optimal strategy need not have a closed form solution.
Figure 3: Do LLM’s tailor response length to problem difficulty? (Left) Average token length produced by GPT-4o on MMLU-Pro Math across prompts in Table 1, split by problems which can be solved without chain-of-thought and problems which NoCoT does not successfully solve. Response lengths are consistently higher for the problems NoCoT unsuccessfully solves. (Right) Token lengths across 500 questions in MMLU-Pro Math for GPT-4o under the DefaultCoT, BeConcise, and WordLimit(10) prompts. Questions are sorted in increasing order of token-complexity across different prompt IDs. Token lengths exhibit a low but non-zero correlation with the true token complexity.
We note that it is unlikely that any feasible prompting technique will achieve the upper bound $\alpha^{*}_{\pi}(T)$ or the lower bound $T^{*}(\alpha)$, as doing so requires (1) knowing the token complexities, (2) inducing a chain-of-thought that exactly matches the token complexity, (3) prioritizing easier questions over harder ones.
Nevertheless $\alpha^{*}_{\pi}(T)$ provides a computable upper bound on maximum accuracy for a particular token budget, which plot in Figure 1 under the label ‘oracle upper bound’, using estimated token complexities $\hat{\tau}_{i}^{\pi}$. Across all the LLMs and benchmarks we consider, we find that this is indeed serves as an upper bound on the performance of the prompting strategies we consider across response lengths, especially for more difficult datasets such as MATH-500 and MMLU-Pro Math. Yet, for GSM8K in Appendix E, we see that this gap with the upper bound $\alpha^{*}_{\pi}(T)$ is much smaller, illustrating that while $\alpha^{*}_{\pi}(T)$ may be challenging to achieve it still gives a reasonable upper bound on performance.
Finally, we circle back to the observation made in Section 2.2 that one can substantially reduce the length of the chain-of-thought (with prompts like BeConcise or NoProperGrammar) while maintaining a similar accuracy to DefaultCoT. This leads to a natural question: what is the lower bound on the number of tokens needed in order to achieve the best possible accuracy? Under the token-complexity hypothesis, we can characterize a simple closed-form expression for this lower bound.
Corollary 1.
Suppose Assumption 1 holds. Let $A^{*}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\{\tau_{i}^{\pi}<\infty\}$ be the maximum possible accuracy achieved by the LLM on the dataset. The number of tokens required to achieve accuracy $\alpha=A^{*}$ under the optimal token allocation:
$T_{\pi}^{*}(A^{*})$ $=\bar{\tau}_{\pi}=\frac{1}{n}\sum_{i=1}^{n}\tau_{i}^{\pi}\mathbf{1}\{\tau_{i}^{\pi}<\infty\}$ (11)
Corollary 1 gives the number of tokens required for ‘lossless compression’, i.e. achieving the best possible accuracy. Remarkably, this lower bound is only the mean of the token-complexities, which can be much smaller than the average token-length of existing prompting strategies. In Table 3, we consider the MMLU-Pro Math dataset across several LLMs, and we compare the lower bound $T_{\pi}^{*}(A^{*})$ with the token counts of DefaultCoT and BeConcise. Across all the LLMs, BeConcise reduces the token counts by 1.2-1.4x, while we see that the theoretical upper bound on token reduction ranges from 3.2-11.2x. Along with the results in Figure 1, this illustrates that while the upper bound may not be exactly attainable, there may be significant room for improvement.
5 Evaluation of Adaptive Compression
Most of all, the token complexity hypothesis emphasizes the importance of adaptive compression: under Assumption 1, improvements in compression only come from matching the token-length $t(X_{i})$ to the token complexity $\tau_{i}^{\pi}$ as closely as possible, rather than from modifications in the chain-of-thought $X_{i}$ itself.
We proceed to show how token complexity provides a framework for studying the capabilities of existing language models to tailor their response length to problem difficulty. Our main observation is that across the chains-of-thought we consider, LLMs do natively adjust response length to the difficulty of the problem, without explicitly being asked to. However, this adaptivity is limited and far from optimal.
Our first evaluation of this capability compares performance of our prompts with a non-adaptive baseline: how does the accuracy-token length trade-off curve compare with a non-adaptive strategy that uses exactly the same number of tokens per question? Note that simply prompting the LLM to use a fixed number of tokens/words/characters/etc., is unable to force a fixed response length Han et al. (2024). To construct this baseline, we let $\bar{\alpha}_{\pi}(T)=\frac{1}{n}\sum\nolimits_{i=1}^{n}\mathbf{1}\{\tau^{\pi}_{i}\leq T\}$ denote the accuracy under the token complexity hypothesis if we use exactly $T$ tokens for each question. Since these values describe the performance under a sub-optimal, non-adaptive strategy, $\bar{\alpha}_{\pi}(T)\leq\alpha^{*}(T)$. The blue line in Figure 1 displays $\bar{\alpha}_{\pi}(T)$ across average token lengths $T$. We observe that across LLMs and benchmarks, the performance of the prompts is close to the non-adaptive baseline $\bar{\alpha}_{\pi}(T)$, with some prompts such as WordLimit(k) or OnlyNumbers exceeding $\bar{\alpha}_{\pi}(T)$. This provides evidence that LLMs outperform a non-adaptive baseline, but not by much.
Model MMLU-Pro Math GSM8K MATH-500 GPT-4o 92.1% 97.0% 89.9% GPT-4o Mini 91.3% 94.1% 91.4% Claude 3.5 Sonnet 92.6% 97.4% 92.1% Claude 3.5 Haiku 90.1% 94.8% 90.2% Llama 3.3 70B 90.6% 96.2% 89.8%
Table 4: Average classification accuracy $\bar{c}_{\pi}$ of threshold classifier $\mathbf{1}\{t(X_{i,k})>\hat{\tau}_{i}^{\pi}\}$ across LLMs and reasoning benchmarks.
Model MMLU-Pro Math GSM8K MATH-500 GPT-4o 5.0% 1.6% 6.0% GPT-4o Mini 5.2% 5.1% 5.7% Claude 3.5 Sonnet 3.4% 1.4% 4.5% Claude 3.5 Haiku 5.1% 4.1% 6.3% Llama 3.3 70B 5.1% 2.7% 6.1%
Table 5: Average relative discrepancy $\text{Err}_{\pi}$ across LLMs and benchmarks. $\text{Err}_{\pi}$ is relative absolute loss between actual performance $\text{Acc}_{\pi}(P_{k})$ and predicted performance $\widehat{\text{Acc}}_{\pi}(P_{k})$, averaged across prompts $P_{k}$.
We can study this in more granularity at the question-level. On the left panel of Figure 3, we consider the average token lengths of GPT-4o on MMLU-Pro Math. We split the problems into two categories: problems that the LLM successfully solves without chain-of-thought (NoCot), and the problems NoCot unsuccessfully solves. Intuitively, the first class of problems are ‘easy’, since they do not require any chain-of-thought, and the rest are harder. Across all the prompts we consider, we observe that the average token-length among ‘easy’ problems is universally smaller than among ‘harder’ problems. Surprisingly, this is even true for the WordLimit/TokenLimit/CharLimit/etc. prompts, which are supposed to limit the LLM’s response to a fixed length. Nevertheless, despite the evidence of adaptive response lengths, this also shows that there is a lot of room for improvement: even though the LLM can solve the problem without any chain-of-thought, the model still proceeds to generate long chains-of-thought.
Through our estimated token complexities, we can perform a much more fine-grained analysis of adaptive response lengths. In the right panel of Figure 3, we look at the token-lengths across questions in MMLU-Pro Math induced by several CoT Prompts – DefaultCoT, BeConcise, and WordLimit(10) – for GPT-4o and Claude 3.5 Sonnet. We sort the questions in increasing order of the estimated token complexities $\tau_{i}^{\pi}$ to assess the degree to which the token lengths produced by the LLM correlate with the true token complexities. We can confirm the observation made earlier: the token lengths exhibit a low but non-zero correlation with the true token complexity. Table 6 in Appendix A measures Kendall-$\tau$ correlation between the token complexities and token lengths across questions for GPT-4o on MMLU-Pro Math, with the highest correlation of 0.53 achieved by OnlyNumbers and the lowest correlation of 0.04 achieved by NoCoT. In total, we highlight that token complexity allows researchers to study the ability to adapt response length to the difficulty of the problem by providing a finer-grained metric of problem difficulty.
6 Towards the Theoretical Limit
The goal of this work is not to propose a new methodology for shortening the response length, and instead our main contribution is to (1) use token complexity to assess existing compression strategies and (2) establish theoretical upper bounds for the optimal tradeoff between reasoning length and performance. Nonetheless, we will briefly discuss how improving adaptivity through predicting problem difficulty can improve upon the universal tradeoff curve observed in in Figure 1 and approach the upper bound $\alpha_{\pi}^{*}(T)$.
Figure 4: Average token length vs Accuracy for Claude 3.5 Sonnet on MMLU-Pro-Math under different prompt routing strategies. Unvertainty-based routing (in purple) does not improve much over existing prompting strategies. However Verifier-based routing (in blue) is able to achieve a better Pareto-frontier, closer to the theoretical upper bound.
As a proof of concept, we consider two prompt-routing strategies, in a similar spirit to Han et al. (2024). First, following the literature on calibration of LLMs Kadavath et al. (2022); Zhu et al. (2023); Chen et al. (2023), we use the language model probabilities in order to use a longer response if the model is uncertain about its current answer. Specifically, the LLM is first prompted with NoCot, i.e. to answer the question without thinking. Then we query the LLM’s token probability of their answer. If the probability is strictly below a threshold, we then query the LLM again with CoT prompt designed to get a longer response (e.g. DefaultCoT). The purple dots in Figure 1 show the performance of this prompt-routing scheme for Claude 3.5 Sonnet on MMLU-Pro Math, where we set the longer prompt to be a selection of four prompts. We observe that the performance is incrementally better but not much so. Moreover, these results are under a very high threshold: we only stick with the answer from NoCoT only if the probability is equal to 1.0. This illustrates that while there are benefits for adapting response length, these benefits may be limited unless one is able to identify question difficulty with high precision.
Nonetheless, if one has access to a perfect verifier, this can substantially improve the compression tradeoff. To evaluate this, we consider another prompt routing strategy, where one instead can verify whether the answer produced by the LLM under NoCoT is correct or not. If the answer is incorrect, then we route to a longer prompt. The blue dots in Figure 4 show performance of verifier-based routing with the same four longer prompts. Unlike uncertainty-based routing, these prompting strategies achieve a significantly better trade-off between reasoning length and performance, and approach the theoretical upper bound. While access to a verifier is unrealistic in practice, this shows how better information of problem difficulty leads to substantial improvements in chain-of-thought compression. This also illustrates that the upper bound $\alpha_{\pi}^{*}(T)$ can be approached and does accurately signify the potential improvements in the accuracy-length tradeoff that can be achieved. Altogether, this demonstrates that our framework sets natural baselines for researchers to compare the performance of length-compressed schemes.
7 Conclusion
Our study presents a systematic investigation into the trade-off between reasoning length and performance in large language models (LLMs), across different prompts. We demonstrate that this trade-off follows a universal Pareto curve, suggesting that reasoning length, rather than specific compression strategies, primarily determines accuracy. Introducing the concept of token complexity, we find that LLM performance at the question-level exhibits a sharp threshold behavior – it is successful only if its chain-of-thought is above a token threshold. Our theoretical analysis, grounded in rate-distortion theory, reveals that existing prompt-based compression strategies operate far from the optimal accuracy-length frontier, highlighting substantial room for improvement. Our work enables researchers to contextualize the performance of new methodologies for improving chain-of-thought compression and assess adaptivity.
Limitations
While our study provides novel insights into the relationship between reasoning length and LLM performance, there are several limitations. First, our study is limited to mathematical reasoning tasks, and it remains to be seen whether similar accuracy-length tradeoffs and token complexity thresholds apply to other domains such as commonsense reasoning, code generation, or open-ended text generation. The theoretical upper bound on accuracy-length tradeoff is rarely achieved due to the practical challenges in estimating token complexities and generating precisely compressed responses. Computing token complexity is computationally expensive, requiring multiple generations per question, making it impractical for large-scale applications. Our approach also assumes that token complexity is well-defined and consistent across different models and tasks, which may not hold for all LLMs or highly complex benchmarks. Furthermore, our experiments were limited to a fixed set of 31 prompts, and exploring a broader range of compression strategies, including model fine-tuning or iterative refinement, could potentially expand the compression frontier. Finally, our analysis is most relevant for strong models on moderately difficult benchmarks; weaker models or extremely challenging tasks may exhibit different behavior.
Appendix A Additional Tables
The following is a table describing the Kendall-$\tau$ correlation between token-lengths with token-complexity across different prompts for GPT-4o on MMLU-Pro Math. While the ordering of which prompts have the highest correlation changes across different models and benchmarks, we observe widely that the correlation ranges between 0 - 0.5.
Prompt ID
Kendall
Correlation
OnlyNumbers 0.53
NoSpaces 0.51
CharLimit(10) 0.50
CharLimit(50) 0.49
NoProperGrammar 0.48
WordLimit(15) 0.25
TokenLimit(10) 0.25
WordLimit(5) 0.20
WordLimit(10) 0.18
NoCoT 0.04
Table 6: GPT-4o on MMLU-Pro Math: Kendall-$\tau$ correlation of token-lengths with token-complexity across different prompts.
Model Prompt Accuracy Token Length GPT-4o NoCoT 41.8 9 DefaultCoT 80.8 585 BeConcise 80.4 415 BulletPoints 75.0 185 OnlyNumbers 79.4 221 NoSpaces 78.8 248 NoProperGrammar 78.8 189 AbbreviateWords 74.2 275 GPT-4o Mini NoCoT 32.4 6 DefaultCoT 75.0 506 BeConcise 74.8 418 BulletPoints 63.0 128 OnlyNumbers 68.8 189 NoSpaces 66.2 219 NoProperGrammar 61.0 119 AbbreviateWords 56.6 259 Claude 3.5 Sonnet NoCoT 48.2 7 DefaultCoT 80.8 324 BeConcise 78.2 227 BulletPoints 75.8 167 OnlyNumbers 76.6 130 NoSpaces 76.8 187 NoProperGrammar 76.8 172 AbbreviateWords 77.2 220 Claude 3.5 Haiku NoCoT 30.2 19 DefaultCoT 68.8 296 BeConcise 68.4 237 BulletPoints 61.8 155 OnlyNumbers 63.2 147 NoSpaces 62.6 193 NoProperGrammar 63.2 169 AbbreviateWords 62.6 243 LLaMA 3.3 70B NoCoT 42.6 20 DefaultCoT 74.4 551 BeConcise 76.6 442 BulletPoints 70.0 193 OnlyNumbers 63.4 304 NoSpaces 67.2 330 NoProperGrammar 68.4 217 AbbreviateWords 70.2 327
Table 7: Accuracy-length tradeoff on MMLU-Pro Math
Model Prompt Accuracy Token Length GPT-4o NoCoT 70.0 30 DefaultCoT 92.8 266 BeConcise 96.6 190 BulletPoints 96.6 104 OnlyNumbers 97.4 76 NoSpaces 96.4 93 NoProperGrammar 95.8 71 AbbreviateWords 95.0 97 GPT-4o Mini NoCoT 28.0 6 DefaultCoT 94.6 292 BeConcise 94.2 216 BulletPoints 92.2 97 OnlyNumbers 92.0 77 NoSpaces 86.6 62 NoProperGrammar 91.0 76 AbbreviateWords 72.8 121 Claude 3.5 Sonnet NoCoT 67.8 7 DefaultCoT 97.0 200 BeConcise 97.4 136 BulletPoints 97.0 100 OnlyNumbers 96.0 66 NoSpaces 97.0 111 NoProperGrammar 97.8 111 AbbreviateWords 95.0 119 Claude 3.5 Haiku NoCoT 30.6 8 DefaultCoT 95.2 211 BeConcise 94.4 169 BulletPoints 92.8 114 OnlyNumbers 93.0 87 NoSpaces 92.4 116 NoProperGrammar 91.8 120 AbbreviateWords 90.8 137 LLaMA 3.3 70B NoCoT 88.6 98 DefaultCoT 96.2 194 BeConcise 95.8 147 BulletPoints 96.2 80 OnlyNumbers 89.6 81 NoSpaces 92.2 102 NoProperGrammar 95.4 91 AbbreviateWords 94.0 136
Table 8: Accuracy-length tradeoff on GSM8K
Model Prompt Accuracy Token Length GPT-4o NoCoT 55.6 116 DefaultCoT 72.8 634 BeConcise 71.6 505 BulletPoints 70.6 302 OnlyNumbers 68.4 272 NoSpaces 71.2 368 NoProperGrammar 68.8 286 AbbreviateWords 72.0 416 GPT-4o Mini NoCoT 25.6 9 DefaultCoT 70.4 610 BeConcise 72.0 528 BulletPoints 68.8 266 OnlyNumbers 66.2 306 NoSpaces 67.6 390 NoProperGrammar 66.8 254 AbbreviateWords 62.0 367 Claude 3.5 Sonnet NoCoT 39.0 9 DefaultCoT 74.8 373 BeConcise 73.0 282 BulletPoints 70.0 203 OnlyNumbers 63.8 172 NoSpaces 70.4 240 NoProperGrammar 69.6 225 AbbreviateWords 71.8 263 Claude 3.5 Haiku NoCoT 25.8 19 DefaultCoT 66.0 373 BeConcise 64.0 286 BulletPoints 60.0 243 OnlyNumbers 59.0 193 NoSpaces 56.4 257 NoProperGrammar 61.4 253 AbbreviateWords 59.4 294 LLaMA 3.3 70B NoCoT 33.8 50 DefaultCoT 55.4 549 BeConcise 67.0 475 BulletPoints 63.6 238 OnlyNumbers 57.0 362 NoSpaces 57.2 439 NoProperGrammar 61.2 294 AbbreviateWords 63.8 383
Table 9: Accuracy-length tradeoff on MATH-500
Appendix B Proof of Theorem 1
Under Assumption 1, we have that the accuracy can be represented as
$\text{Acc}_{\pi}(P)=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\{\tau_{i}\geq t_{i}\}$ (12)
where we suppress dependence on $\pi$ for now. The optimization problem then becomes
$\alpha^{*}_{\pi}(T)=\max_{t}$ $\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\{\tau_{i}\geq t_{i}\}$ (13)
s.t. $\frac{1}{n}\sum_{i=1}^{n}t_{i}\leq T$ (14)
Note that the optimal strategy is either to set $t_{i}=\tau_{i}$ or 00. Thus, this is equivalent to a Knapsack problem,
$\alpha^{*}_{\pi}(T)=\max_{x_{i}\in\{0,1\}}$ $\frac{1}{n}\sum_{i=1}^{n}x_{i}$ (15)
s.t. $\frac{1}{n}\sum_{i=1}^{n}\tau_{i}x_{i}\leq T$ (16)
where $x_{i}$ are indicators whether the LLM tries to solve the problem or not. This is a special case where there are unit rewards (since all problems are weighted equally). The optimal strategy in this case is a greedy policy, sorting the questions in increasing order of token complexity $\tau_{(1)}<...<\tau_{(n)}$ and only solving as many as can fit within budget. $t_{T}\equiv\inf\{t\in\mathbb{R}:E_{n}(t)=t\}$ is the highest value of $\tau$ that the LLM puts into the knapsack after ordering them greedily, so thus $\alpha_{n}^{*}(T)$ is the total number of questions with token complexity less than $\tau$. The proof for $T^{*}(\alpha)$ proceeds similarly, as the optimal strategy is identical.
Appendix C Example Prompt
We use the following template for our prompts:
Answer the following question. {PROMPT}
Question: {QUESTION}
The last line of your response
should be of the following format:
Answer: \$ANSWER (without quotes)
where ANSWER is your final answer.
Appendix D Tradeoff curves for more models and benchmarks
In this section, we present results for the performance of different models and benchmarks (GSM8K and MATH-500). We see broadly that performance across all prompts lies on the same trade-off curve.
Figure 5: Anthropic Claude 3.5 Performance on GSM8K
Figure 6: OpenAI GPT-4o Performance on GSM8K
Figure 7: Meta Llama 3.3-70B Performance on GSM8K
Figure 8: Anthropic Claude 3.5 Haiku Performance on GSM8K
Figure 9: OpenAI GPT-4o Mini Performance on GSM8K
Figure 10: Anthropic Claude 3.5 Performance on MATH-500
Figure 11: OpenAI GPT-4o Performance on MATH-500
Figure 12: Meta Llama 3.3-70B Performance on MATH-500
Figure 13: Anthropic Claude 3.5 Haiku Performance on MATH-500
Figure 14: OpenAI GPT-4o-Mini Performance on MATH-500
Appendix E Actual accuracy vs Predicted Accuracy from Token Complexity
Figure 15: Anthropic Claude 3.5 Prediction on GSM8K
Figure 16: OpenAI GPT-4o Prediction on GSM8K
Figure 17: Meta Llama 3.3-70B Prediction on GSM8K
Figure 18: OpenAI GPT-4o-Mini Prediction on GSM8K
Figure 19: Anthropic Claude 3.5 Haiku Prediction on GSM8K
Figure 20: Anthropic Claude 3.5 Prediction on MATH-500
Figure 21: OpenAI GPT-4o Prediction on MATH-500
Figure 22: Meta Llama 3.3-70B Prediction on MATH-500
Figure 23: Anthropic Claude 3.5 Haiku Prediction on MATH-500
Figure 24: GPT-4o-Mini Prediction on MATH-500